# Optimal Deterministic Rendezvous in Labeled Lines

### v1

@@ 2.1 Notation
Throughout the paper, $\log$ refers to the base-$2$ logarithm, i.e., $\log(2^{x})=x$ for any $x\in\mathbb{R}$. For each $x\in\mathbb{R}$, the $\log^{*}$ function is defined as $\log^{*}(x)=1$ if $x\leq 1$, $\log^{*}(x)=1+\log^{*}(\log(x))$. This is off-by-one from the maybe more common definition that sets $\log^{*}(x)=0$ if $x\leq 1$. We do so to have that $\Theta(D\log^{*}\ell)$ is of order $\Theta(D)$ rather than 00 when $\ell=1$. For non-zero real $a$ and positive integer $b$, we define $a\mathbin{\uparrow\uparrow}0=1$ and $a\mathbin{\uparrow\uparrow}b=a^{a\mathbin{\uparrow\uparrow}(b-1)}$. For any positive integer $k$ and real $x$ s.t. $2\mathbin{\uparrow\uparrow}(k-1)<x\leq 2\mathbin{\uparrow\uparrow}k$, $\log^{*}(x)=k+1$.
Graph notation
We write $G=(V,E)$ for a graph over vertices $V$ and edges $E$. Each node $v\in V$ is equipped with a unique label or identifier $\ell(v)\in\mathcal{L}$. We identify the set of labels with the positive integers, $\mathcal{L}=\mathbb{N}^{*}$. We partition the nodes according to their labels and the $\log^{*}$ function, defining $V_{i}=\{v\in V:\log^{*}(\ell(v))=i\}$, so that $V=\bigsqcup_{i\geq 1}V_{i}$. We also define $V_{\leq i}=\bigsqcup_{j=1}^{i}V_{j}$ and $V_{<i}=V_{\leq i}\setminus V_{i}$. The degree of a node $v\in V$ is denoted by $\deg(v)$, the distance $\operatorname{dist}(v,v^{\prime})$ between two nodes corresponds to the (edge-)length of a shortest path between them.
@@
The Labeled Infinite Line
@@ 2.2 Computational Models
Our Mobile Agent Model
The agents are placed and move in a network modeled by a graph $G=(V,E)$. $G$ can be finite or countably infinite. Each node $v\in V$ of the graph has $\deg(v)$ ports, and each edge is the graph is matched to exactly one port of each of its two endpoints. The ports of a node $v$ are labeled 00 to $\deg(v)-1$. At a node $v$, an agent can read the node’s label $\ell(v)$ if the graph is labeled, see the node’s degree $\deg(v)$, and distinguish the $\deg(v)$ ports leading out of the node.
The agents execute the same deterministic algorithm, as in, when put in the same configuration, agents $\alpha$ and $\beta$ perform the same set of moves. Time is divided into discrete rounds. The agents are initially asleep, and upon wake-up (or startup), they have no notion of how much time has elapsed before. Formally, the agents have an internal clock that is initialized at 00 upon waking-up, and from there on keeps track of how many rounds passed since wake-up. We denote by $\tau$ the delay between the wake-up times of the two agents. The agents can be thought of as waking up in rounds 00 and $\tau$. The initial distance between the two agents is denoted by $D$.
@@
The LOCAL Model on Infinite Graphs
@@
In infinite graphs, it is nonsensical to consider the complexity of most problems as the time needed for the last node to terminate. The complexity also cannot be measured with respect to the number of nodes or the value of the largest identifier. To tame the infinite, we instead aim to bound the time it takes a node to finish with respect to some local parameters, for example, its identifier. To have nodes terminate at different times, the problems we consider must be inherently quite local. We refer to this model, where nodes commit to an output in a time that is a function of a local parameter, and thus algorithms make sense on both finite and infinite graphs, as the EarlyStop-LOCAL model (ES-LOCAL for short). This model was also used in the work by Miller and Pelc [17] and appeared previously in other sources [10, Exercise 1.7].
@@ 2.3 Useful Subroutines
Let each node $v\in V$ be given a list of colors $\psi(v)$. A list-coloring of the graph for these lists is a map $\chi:V\to\bigcup_{v\in V}\psi(v)$ s.t. for each node $v$, $\chi(v)\in\psi(v)$, and for any two adjacent nodes $u$ and $v$, $\chi(u)\neq\chi(v)$. When each node $v$ is given a list of size $\lvert\psi(v)\rvert\geq\deg(v)+1$, the coloring is known as a degree+1-list coloring. Having lists with this minimum size guarantees the existence of a solution $\chi$. In fact, it also guarantees that a partial solution can always be extended to a full one. When the lists are all equal to $[k]$ for some integer $k$, the problem is known as $k$-coloring. In the extensively studied $\Delta+1$-coloring problem, each node has $[\Delta+1]$ as its list of colors, where $\Delta$ is an upper bound on the maximum degree of the graph.
@@ 3.1 Ensuring Rendezvous from Crossing: CarefulWalk
We give a simple procedure (Algorithm 1) to cross edges which ensures that the agents cannot exchange positions on the line without achieving rendezvous. To do so, CarefulWalk takes $4$ rounds to cross each edge. Its basic idea is to have the agents behave differently depending on whether they cross an edge in the direction of increasing or decreasing IDs. The same can be achieved given a common orientation of the edges instead of node labels.
@@
2:Wait at initial node for one round.$\triangleright$ Agent at $u$ at $t=0,1$.
@@
5:     Wait two rounds. $\triangleright$ Agent at $u$ at $t=2,3,4$.
6$\triangleright$ $\operatorname{ID}(v)<\operatorname{ID}(u)$
@@
Figure 1: Illustration of the moves performed by CarefulWalk to cross an edge $uv$. The five circles below each 00 (resp. $1$) represent the node of lower ID (resp. higher ID) at $5$ successive time steps. Arrows represent the moves (or absence thereof) of an agent in a CarefulWalk crossing the edge. The left and middle figures depict crossing the edge in opposite directions. The right figure depicts two agents performing a CarefulWalk over the same edge simultaneously in opposite directions, achieving rendezvous at time step $4$, after $3$ moves. The $5$-bit strings in the figures describe the positions occupied the agents over the $5$ time steps.
Consider an algorithm $\mathcal{A}$ for the moves of an agent. Let $\operatorname{care}(\mathcal{A})$ be the algorithm obtained from $\mathcal{A}$ by replacing each edge-crossing in $\mathcal{A}$ by a call to CarefulWalk, and stretching each stay at a node by a factor of $4$. This new algorithm basically does the same moves as $\mathcal{A}$ at a slower pace. We show in Theorem 3.1 that this transformed algorithm is able to achieve rendezvous as $\mathcal{A}$ would in the stronger setting where the agents can detect simultaneous edge-crossing.
@@
Let agent $\alpha$ be the first agent to wake up (in round 00), and $\beta$ the other agent (waking up in round $\tau\geq 0$). Consider the execution of $\operatorname{care}(\mathcal{A})$. In that execution, for each integer $i\geq 0$, let $x(i)$ be the position of agent $\alpha$ in round $i$, and $y(i)$ that of agent $\beta$. In particular, $y(i)=y(0)$ for all $i\leq\tau$.
@@
Suppose that $x^{\prime}({t^{*}})=y^{\prime}({t^{*}})$. If $r=0$ (i.e., $\tau=0\mod 4$), then $\alpha$ and $\beta$ are at the same node $x(4{t^{*}})=y(4{t^{*}})$ in round $4{t^{*}}$. Otherwise ($r\in\{1,2,3\}$), consider the positions of the agents in rounds $4{t^{*}}$ and $24{t^{*}}-$ 2. Regardless of whether $\alpha$ was staying idle for $4$ rounds or performing a CarefulWalk to node $x(4{t^{*}})$ in the $4$ rounds prior to round $4{t^{*}}$, it was at node $x(4{t^{*}})$ in rounds $4{t^{*}}$ and $24{t^{*}}-$ 2 (see Algorithm 1 and Fig. 1, the description and an illustration of CarefulWalk). Similarly, regardless of the action that agent $\beta$ initiates in round $4{t^{*}}-r$, it is at node $y(4{t^{*}}-r)=x(4{t^{*}})$ in rounds $(4{t^{*}}-r)$ and $(4{t^{*}}-r+1)$. For all possible values of $r$, the agents meet in either round $4{t^{*}}$ or $24{t^{*}}-$ 2.
Suppose that $x^{\prime}({t^{*}})=y^{\prime}({t^{*}}-1)$ and $x^{\prime}({t^{*}}-1)=y^{\prime}({t^{*}})$. A similar case analysis as the one we just did shows that the agents necessarily meet in a round in the interval $[4{t^{*}}-4,4{t^{*}}-1]$ (see Fig. 2).
@@
Theorem 3.1 allows us to design algorithms for the simpler setting where agents detect simultaneous edge-crossing, and automatically translate rendezvous results about these algorithms to the setting where agents only achieve rendezvous if at the same node in a given round, which is the original setting of interest. This dramatically simplifies the work of the algorithm designer. Importantly, on the line, having a detection mechanism for simultaneous edge-crossing prevents the agents to swap positions without achieving rendezvous.
@@ 3.2 Computing Ruling Sets on Paths
8:for $6$ iterations do
9:     For each $v\in U$, compute $b_{v}=\min_{u\in S_{\mathrm{out}}}\operatorname{dist}(v,u)$. Let $U_{\geq R}=\{v\in U:b_{v}\geq R\}$.
10:     Add to $S_{\mathrm{out}}$ all $v\in U_{\geq R}$ s.t. $\forall u\in N^{R-1}(v)\cap U_{\geq R},(b_{u}<b_{v}\vee(b_{u}=b_{v}\wedge\ell(v)>\ell(u)))$.
@@
Algorithm 2 computes a $U$-limited $(R,R-1)$ ruling set of $G$ in $O(R\log^{*}n)$ rounds of deterministic LOCAL.
@@
$d=\lceil\log R\rceil$ implies that $2^{d}\geq R>2^{d-1}$. We show that the following properties always hold for each integer $i\in[0,d-1]$: the nodes in the set $S_{i}$ are at distance at least $2^{i}$ from each other (packing property), and every node from $U$ is at distance at most $22^{i+1}-i-$ 2 from a node in $S_{i}$ (covering property). We then analyze the distances for the last two sets $S_{d}$ and $S_{\mathrm{out}}$. Finally, we use those properties to guarantee the $O(R\log^{*}n)$ runtime.
@@
For $i=0$, the property trivially holds: $12^{i}=$ 1 and two distinct nodes are necessarily at distance at least $1$ from each other. For $i\in[1,d-1]$, the property follows directly from the fact that in iteration $i$ of the loop, we compute a maximal independent set over $H_{i}=G^{R_{i}}[S_{i-1}]$. Since the nodes in $S_{i}$ form an independent set in $G^{R_{i}}$, any two nodes in $S_{i}$ are not adjacent in $G^{R_{i}}$, thus are at distance at least $R_{i}+1=2^{i}$ from each other in $G$, giving the claim. Additionally, nodes in $S_{d}$ are at least $R_{d}+1=R$-apart by the same argument. Finally, the nodes that are added to $S_{d}$ to obtain $S_{\mathrm{out}}$ are selected to be at distance $\geq R$ from other nodes already in $S_{\mathrm{out}}$ (condition $b_{v}\geq R$ in computing $U_{\geq R}$), and so that two nodes at distance $<R$ are never simultaneously added to $S_{\mathrm{out}}$ (condition that $b_{v}$ is maximal among $U_{\geq R}$ nodes within a $(R-1)$-radius, using IDs as tie-breakers)
@@
Since $S_{0}=U$, for $i=0$, we have that every node in $U$ is at distance at most $20=2^{i+1}-i-$ 2 from a node in $S_{i}$.
@@
Therefore for every integer $i\in[1,d-1]$, the maximum distance of a node in $U$ to a node in $S_{i}$ is at most $\sum_{j=1}^{i}R_{j}=\sum_{j=1}^{i}(2^{j}-1)=2^{i+1}-i-2\leq 2^{d}-d-1\leq 2R-2$. The MIS computation over $S_{d}$ adds at most $R-1$ to this distance, giving a maximum distance of $33R-$ 3.
Finally, consider the computation of $S_{\mathrm{out}}$, initialized as $S_{d}$. Consider an iteration of the second loop, and take the current values of $S_{\mathrm{out}}$ and $(b_{v})_{v\in U}$. Consider a node $v\in U$ s.t. $b_{v}\geq R$. A node in its distance-$(3R-4)$ neighborhood must be added to $S_{\mathrm{out}}$. Indeed, we have $b_{u}\leq 3R-3$ for all nodes $u\in U$ from the established properties of $S_{d}$. This implies that within distance $32R-$ 3 from $v$, there is a node $w$ s.t. $b_{w}\geq R$ and $b_{w}\geq b_{u}$ for all $u\in N^{R-1}(w)$. Either this node is added to $S_{\mathrm{out}}$, or a node $u\in N^{R-1}(w)$ of equal value $b_{u}=b_{w}$ is added to $S_{\mathrm{out}}$.
Suppose that after $k$ iterations of the loop, for some node $v\in U$, it holds that $b_{v}\geq R$. This means that we had $b_{v}\geq R$ for the $k$ iterations of the loop, and therefore, $k$ nodes were added to $S_{\mathrm{out}}$ in the distance-$(3R-4)$ neighborhood of $v$. A subsequent iteration number $k+1$ should add another node to $S_{\mathrm{out}}$ within this neighborhood. But nodes in $S_{\mathrm{out}}$ are at distance at least $R$ from one another, so the distance-$(3R-4)$ neighborhood of $v$ can contain at most $6$ nodes in $S_{\mathrm{out}}$. Therefore, after $6$ iterations of the loop, all nodes $v\in U$ satisfy $b_{v}\leq R-1$, i.e., are at distance $<R$ from $S_{\mathrm{out}}$.
@@
$G$ being a path/line, the fact that two nodes in $S_{i-1}$ are at distance at least $22^{i-1}\geq(R_{i}+1)/$ 2 from one another in $G$ bounds the maximum degree of $H_{i}=G^{R_{i}}[S_{i-1}]$ to $2$. The nodes of these graphs are also all from $U$, which bounds their identifiers to be at most $n$. These two bounds imply that all the computations we do on each of these graphs can be done in $O(\log^{*}n)$ rounds of LOCAL on the said graph. Simulating a round of LOCAL on a graph $H\subseteq G^{k}$ can be done in $k$ rounds of communication over $G$. As we consider $\Theta(\log R)$ graphs whose costs of simulation increase geometrically, we get a total complexity of $O(\sum_{i=1}^{d}R_{i}\log^{*}n)=O(R\log^{*}n)$ for the $d$ MIS computations. The greedy extension of $S_{d}$ to $S_{\mathrm{out}}$ only adds $O(R)$ rounds to this complexity, as $O(R)$ rounds suffice for each node to compute its distance $\leq 3R-3$ to the nearest node in $S_{d}$ and compare it to the values of the nodes within distance $R-1$. ∎

### v2

@@ 2.1 Notation
Throughout the paper, $\log$ refers to the base-$2$ logarithm, i.e., $\log(2^{x})=x$ for any $x\in\mathbb{R}$. For each $x\in\mathbb{R}$, the $\log^{*}$ function is defined as $\log^{*}(x)=1$ if $x\leq 1$, $\log^{*}(x)=1+\log^{*}(\log(x))$. This is off-by-one from the maybe more common definition that sets $\log^{*}(x)=0$ if $x\leq 1$. We do so to have that $\Theta(D\log^{*}\ell)$ is of order $\Theta(D)$ rather than 0 when $\ell=1$.
Graph notation.
We write $G=(V,E)$ for a graph over vertices $V$ and edges $E$. Each node $v\in V$ is equipped with a unique label or identifier $\ell(v)\in\mathcal{L}$. We identify the set of labels with the positive integers, $\mathcal{L}=\mathbb{N}^{*}$. We partition the nodes according to their labels and the $\log^{*}$ function, defining $V_{i}=\{v\in V:\log^{*}(\ell(v))=i\}$, so that $V=\bigsqcup_{i\geq 1}V_{i}$. We also define $V_{\leq i}=\bigsqcup_{j=1}^{i}V_{j}$ and $V_{<i}=V_{\leq i}\setminus V_{i}$. The degree of a node $v\in V$ is denoted by $\deg(v)$, the distance $\operatorname{dist}(v,v^{\prime})$ between two nodes corresponds to the (edge-)length of a shortest path between them, and $N(v)$ is the set of neighbors of $v$. For every integer $k$, $N^{k}(v)=\{u\in V:\operatorname{dist}(u,v)\in[1,k]\}$ is the $k$-hop neighborhood of $v$.
@@
The Labeled Infinite Line.
@@ 2.2 Computational Models
Our Mobile Agent Model.
The agents are placed and move in a network modeled by a graph $G=(V,E)$. $G$ can be finite or countably infinite. Each node $v\in V$ of the graph has $\deg(v)$ ports, and each edge is the graph is matched to exactly one port of each of its two endpoints. The ports of a node $v$ are labeled 0 to $\deg(v)-1$. At a node $v$, an agent can read the node’s label $\ell(v)$ if the graph is labeled, see the node’s degree $\deg(v)$, and distinguish the $\deg(v)$ ports leading out of the node.
The agents execute the same deterministic algorithm, as in, when put in the same configuration, agents $\alpha$ and $\beta$ perform the same set of moves. Time is divided into discrete rounds. The agents are initially asleep, and upon wake-up (or startup), they have no notion of how much time has elapsed before. Formally, the agents have an internal clock that is initialized at 0 upon waking-up, and from there on keeps track of how many rounds passed since wake-up. We denote by $\tau$ the delay between the wake-up times of the two agents. The agents can be thought of as waking up in rounds 0 and $\tau$. The initial distance between the two agents is denoted by $D$.
@@
The LOCAL Model on Infinite Graphs.
@@
In infinite graphs, it is nonsensical to consider the complexity of most problems as the time needed for the last node to terminate. The complexity also cannot be measured with respect to the number of nodes or the value of the largest identifier. To tame the infinite, we instead aim to bound the time it takes a node to finish with respect to some local parameters, for example, its identifier. To have nodes terminate at different times, the problems we consider must be inherently quite local. We refer to this model, where nodes commit to an output in a time that is a function of a local parameter, and thus algorithms make sense on both finite and infinite graphs, as the EarlyStop-LOCAL model (ES-LOCAL for short). This model was also used in the work by Miller and Pelc [MP25] and appeared previously in other sources [HS20, Exercise 1.7].
@@ 2.3 Useful Subroutines
Let each node $v\in V$ be given a list of colors $\psi(v)$. A list-coloring of the graph for these lists is a map $\chi:V\to\bigcup_{v\in V}\psi(v)$ s.t. for each node $v$, $\chi(v)\in\psi(v)$, and for any two adjacent nodes $u$ and $v$, $\chi(u)\neq\chi(v)$. When each node $v$ is given a list of size $\lvert\psi(v)\rvert\geq\deg(v)+1$, the coloring is known as a degree+1-list coloring. Having lists with this minimum size guarantees the existence of a solution $\chi$. In fact, it also guarantees that a partial solution can always be extended to a full one. When the lists are all equal to $[k]$ for some integer $k$, the problem is known as $k$-coloring. In the extensively studied $\Delta+1$-coloring problem, each node has $\Delta+1]$ as its list of colors, where $\Delta$ is an upper bound on the maximum degree of the graph.
@@ 3.1 Ensuring Rendezvous from Crossing: CarefulWalk
We give a simple procedure (Algorithm˜1) to cross edges which ensures that the agents cannot exchange positions on the line without achieving rendezvous. To do so, CarefulWalk takes $4$ rounds to cross each edge. Its basic idea is to have the agents behave differently depending on whether they cross an edge in the direction of increasing or decreasing IDs. The same can be achieved given a common orientation of the edges instead of node labels.
@@
[…]
@@
5:  Wait two rounds. $\triangleright$ Agent at $v$ at $t=2,3,4$.
[…]
@@
Figure 1: Illustration of the moves performed by CarefulWalk to cross an edge uvuv. The five circles below each 0 (resp. $1$) represent the node of lower ID (resp. higher ID) at 55 successive time steps. Arrows represent the moves (or absence thereof) of an agent in a CarefulWalk crossing the edge. The left and middle figures depict crossing the edge in opposite directions. The right figure depicts two agents performing a CarefulWalk over the same edge simultaneously in opposite directions, achieving rendezvous at time step 44, after 33 moves. The 55-bit strings in the figures describe the positions occupied the agents over the $5$ time steps.
Consider an algorithm $\mathcal{A}$ for the moves of an agent. Let $\operatorname{care}(\mathcal{A})$ be the algorithm obtained from $\mathcal{A}$ by replacing each edge-crossing in $\mathcal{A}$ by a call to CarefulWalk, and stretching each stay at a node by a factor of $4$. This new algorithm basically does the same moves as $\mathcal{A}$ at a slower pace. We show in Theorem˜3.1 that this transformed algorithm is able to achieve rendezvous as $\mathcal{A}$ would in the stronger setting where the agents can detect simultaneous edge-crossing.
@@
Let agent $\alpha$ be the first agent to wake up (in round 0), and $\beta$ the other agent (waking up in round $\tau\geq 0$). Consider the execution of $\operatorname{care}(\mathcal{A})$. In that execution, for each integer $i\geq 0$, let $x(i)$ be the position of agent $\alpha$ in round $i$, and $y(i)$ that of agent $\beta$. In particular, $y(i)=y(0)$ for all $i\leq\tau$.
@@
Suppose that $x^{\prime}({t^{*}})=y^{\prime}({t^{*}})$. If $r=0$ (i.e., $\tau=0\mod 4$), then $\alpha$ and $\beta$ are at the same node $x(4{t^{*}})=y(4{t^{*}})$ in round $4{t^{*}}$. Otherwise ($r\in\{1,2,3\}$), consider the positions of the agents in rounds $4{t^{*}}$ and $4{t^{*}}-2$. Regardless of whether $\alpha$ was staying idle for $4$ rounds or performing a CarefulWalk to node $x(4{t^{*}})$ in the $4$ rounds prior to round $4{t^{*}}$, it was at node $x(4{t^{*}})$ in rounds $4{t^{*}}$ and $4{t^{*}}-2$ (see Algorithm˜1 and Fig.˜1, the description and an illustration of CarefulWalk). Similarly, regardless of the action that agent $\beta$ initiates in round $4{t^{*}}-r$, it is at node $y(4{t^{*}}-r)=x(4{t^{*}})$ in rounds $(4{t^{*}}-r)$ and $(4{t^{*}}-r+1)$. For all possible values of $r$, the agents meet in either round $4{t^{*}}$ or $4{t^{*}}-2$.
Suppose that $x^{\prime}({t^{*}})=y^{\prime}({t^{*}}-1)$ and $x^{\prime}({t^{*}}-1)=y^{\prime}({t^{*}})$. A similar case analysis as the one we just did shows that the agents necessarily meet in a round in the interval $4{t^{*}}-4,4{t^{*}}-1]$ (see Fig.˜2).
@@
Theorem˜3.1 allows us to design algorithms for the simpler setting where agents detect simultaneous edge-crossing, and automatically translate rendezvous results about these algorithms to the setting where agents only achieve rendezvous if at the same node in a given round, which is the original setting of interest. This dramatically simplifies the work of the algorithm designer. Importantly, on the line, having a detection mechanism for simultaneous edge-crossing prevents the agents to swap positions without achieving rendezvous.
@@ 3.2 Computing Ruling Sets on Paths
8:for 66 iterations do
9:  For each $v\in U$, compute $b_{v}=\min_{u\in S_{\mathrm{out}}}\operatorname{dist}(v,u)$. Let $U^{\geq R}=\{v\in U:b_{v}\geq R\}$.
10:  Add to $S_{\mathrm{out}}$ all $v\in U^{\geq R}$ s.t. $\forall u\in N^{R-1}(v)\cap U^{\geq R},(b_{u}<b_{v}\vee(b_{u}=b_{v}\wedge\ell(v)>\ell(u)))$.
@@
Executed by a set of nodes $U$ with identifiers between $1$ and $n$ on an infinite labeled line, Algorithm˜2 computes a $U$-limited (R,R1)(R,R-1) ruling set of $G$ in $O(R\log^{*}n)$ rounds of deterministic LOCAL.
@@
$d=\lceil\log R\rceil$ implies that $2^{d}\geq R>2^{d-1}$. We show that the following properties always hold for each integer $i\in[0,d-1]$: the nodes in the set $S_{i}$ are at distance at least $2^{i}$ from each other (packing property), and every node from $U$ is at distance at most $2^{i+1}-i-2$ from a node in $S_{i}$ (covering property). We then analyze the distances for the last two sets $S_{d}$ and $S_{\mathrm{out}}$. Finally, we use those properties to guarantee the $O(R\log^{*}n)$ runtime.
@@
For $i=0$, the property trivially holds: $2^{i}=1$ and two distinct nodes are necessarily at distance at least $1$ from each other. For $i\in[1,d-1]$, the property follows directly from the fact that in iteration $i$ of the loop, we compute a maximal independent set over $H_{i}=G^{R_{i}}[S_{i-1}]$. Since the nodes in $S_{i}$ form an independent set in $G^{R_{i}}$, any two nodes in $S_{i}$ are not adjacent in $G^{R_{i}}$, thus are at distance at least $R_{i}+1=2^{i}$ from each other in $G$, giving the claim. Additionally, nodes in $S_{d}$ are at least $R_{d}+1=R$-apart by the same argument. Finally, the nodes that are added to $S_{d}$ to obtain $S_{\mathrm{out}}$ are selected to be at distance $\geq R$ from other nodes already in $S_{\mathrm{out}}$ (condition $b_{v}\geq R$ in computing $U^{\geq R}$), and so that two nodes at distance <R<R are never simultaneously added to $S_{\mathrm{out}}$ (condition that $b_{v}$ is maximal among $U^{\geq R}$ nodes within a (R1)(R-1)-radius, using IDs as tie-breakers)
@@
Since $S_{0}=U$, for $i=0$, we have that every node in $U$ is at distance at most $0=2^{i+1}-i-2$ from a node in $S_{i}$.
@@
Therefore for every integer $i\in[1,d-1]$, the maximum distance of a node in $U$ to a node in $S_{i}$ is at most $\sum_{j=1}^{i}R_{j}=\sum_{j=1}^{i}(2^{j}-1)=2^{i+1}-i-2\leq 2^{d}-d-1\leq 2R-2$. The MIS computation over $S_{d}$ adds at most R1R-1 to this distance, giving a maximum distance of 3R33R-3.
Finally, consider the computation of $S_{\mathrm{out}}$, initialized as $S_{d}$. Consider an iteration of the second loop, and take the current values of $S_{\mathrm{out}}$ and $(b_{v})_{v\in U}$. Consider a node $v\in U$ s.t. $b_{v}\geq R$. A node in its distance-(3R4)(3R-4) neighborhood must be added to $S_{\mathrm{out}}$. Indeed, we have $b_{u}\leq 3R-3$ for all nodes $u\in U$ from the established properties of $S_{d}$. This implies that within distance 2R32R-3 from $v$, there is a node $w$ s.t. $b_{w}\geq R$ and $b_{w}\geq b_{u}$ for all $u\in N^{R-1}(w)$. Either this node is added to $S_{\mathrm{out}}$, or a node $u\in N^{R-1}(w)$ of equal value $b_{u}=b_{w}$ is added to $S_{\mathrm{out}}$.
Suppose that after $k$ iterations of the loop, for some node $v\in U$, it holds that $b_{v}\geq R$. This means that we had $b_{v}\geq R$ for the $k$ iterations of the loop, and therefore, $k$ nodes were added to $S_{\mathrm{out}}$ in the distance-(3R4)(3R-4) neighborhood of $v$. A subsequent iteration number $k+1$ should add another node to $S_{\mathrm{out}}$ within this neighborhood. But nodes in $S_{\mathrm{out}}$ are at distance at least $R$ from one another, so the distance-(3R4)(3R-4) neighborhood of $v$ can contain at most $6$ nodes in $S_{\mathrm{out}}$. Therefore, after $6$ iterations of the loop, all nodes $v\in U$ satisfy $b_{v}\leq R-1$, i.e., are at distance <R<R from $S_{\mathrm{out}}$.
@@
$G$ being a path/line, the fact that two nodes in $S_{i-1}$ are at distance at least $2^{i-1}\geq(R_{i}+1)/2$ from one another in $G$ bounds the maximum degree of $H_{i}=G^{R_{i}}[S_{i-1}]$ to $2$. The nodes of these graphs are also all from $U$, which bounds their identifiers to be at most $n$. These two bounds imply that all the computations we do on each of these graphs can be done in $O(\log^{*}n)$ rounds of LOCAL on the said graph. Simulating a round of LOCAL on a graph $H\subseteq G^{k}$ can be done in $k$ rounds of communication over $G$. As we consider $\Theta(\log R)$ graphs whose costs of simulation increase geometrically, we get a total complexity of $O(\sum_{i=1}^{d}R_{i}\log^{*}n)=O(R\log^{*}n)$ for the $d$ MIS computations. The greedy extension of $S_{d}$ to $S_{\mathrm{out}}$ only adds $O(R)$ rounds to this complexity, as $O(R)$ rounds suffice for each node to compute its distance $\leq 3R-3$ to the nearest node in $S_{d}$ and compare it to the values of the nodes within distance R1R-1. ∎